OpenStudy (anonymous):

what is the equation of the line slope 5 and y-intercept-2?

OpenStudy (ccswims):

y = 5x - 2

OpenStudy (ccswims):

do you understand or do you want me to explain this?

OpenStudy (anonymous):

please explain my choices are a. 5x+y=2 b. 2x-y=5 c. 5x-y=2 d. 2x+y-5

OpenStudy (ccswims):

so they want it in standard form

OpenStudy (ccswims):

it would be C

OpenStudy (ccswims):

y = 5x - 2 x needs to be in the beginning so cancel out y 5x - y = 2

OpenStudy (ccswims):

Standard form Ax - By = C

OpenStudy (anonymous):

So, when you do this, you need to change the sign of 5x because its going over the equal sign. Corret? so over all once you get, y=mx+-B you just subtract y.

OpenStudy (ccswims):

yes

OpenStudy (ccswims):

and this cause -2 to become 2

OpenStudy (anonymous):

The equal sign always needs to go at the end? Or is this just because of the answer choices

OpenStudy (ccswims):

always, it's just like a two-step equation

OpenStudy (anonymous):

Can I ask you to do another question? Explaining it step by step?

OpenStudy (anonymous):

Which is an equation of the line with slope -2/3 passing through the point (-10,0)?

OpenStudy (ccswims):

ok so that equation would start like y = mx + b

OpenStudy (ccswims):

m is your slope. You know your slope is -2/3 so plug that in

OpenStudy (anonymous):

Yes, I just have trouble with the B

OpenStudy (ccswims):

b = 0 so just put nothing, (-10,0) the y - intercept = 0

OpenStudy (anonymous):

So, the Y in your coordinate is always B?

OpenStudy (ccswims):

yes b = y-intercept

OpenStudy (anonymous):

Ok, what would be the answer? If i plug it in, it's not one of my answer choices

OpenStudy (ccswims):

what are the answer choices

OpenStudy (anonymous):

A. Y= -2/3x -10 B. 2x+3y = 10 C. Y= +10 = -2x D. Y= -2/3x - 20/3

OpenStudy (ccswims):

B

OpenStudy (anonymous):

Ok, so the numerator is always x?

OpenStudy (ccswims):

No, here wait let me explain more

OpenStudy (ccswims):

in standard form x always has to be positive and it has to be rational. So -2/3x + y = 10 is unacceptable you must cancel out the fraction by doing 3/1 times -2/3, canceling out the 3 and leaving the 2 behind What you do to x you must do to y So y times 3 is 3y

OpenStudy (ccswims):

Leaving it 2x + 3y = 10

OpenStudy (anonymous):

-2/3x + y =10?

OpenStudy (ccswims):

yes, standard form is Ax + By = C Our first equation was y = -2/3x x must be in the beginning to move the -2/3x to the front and add + y

OpenStudy (anonymous):

I'm only in 8th grade, I have no idea what this all means, I understand the Standard Form, but "x must be in the beginning to move the -2/3x to the front and add + y" I have no clue on.

OpenStudy (ccswims):

So you understand Standard Form, you know which ever number (rational or irrational) that has the variable "x" must be in the beginning

OpenStudy (anonymous):

Standard form = No Fractions or Mixed number's, Correct?

OpenStudy (ccswims):

Yes, and which ever number holds the variable "x" it MUST be positive

OpenStudy (ccswims):

So "x" must be rational and positive

OpenStudy (ccswims):

"y" doesn't have too

OpenStudy (anonymous):

So, Y= -2/3, would be 2x+3y = 0? Because 0 is the Y Intercept

OpenStudy (ccswims):

no the constant is on the other side of the equal sign, not the y - intercept

OpenStudy (anonymous):

Constant? -1 - (-100) and 1-100? All even and positive number's? That probably something stupid to know, its just tomorrow I have a test on this but without the answer choices.

OpenStudy (anonymous):

to ask, not to know

OpenStudy (ccswims):

a constant is any number that does not hold a variable

OpenStudy (ccswims):

2x is not a constant but -46 is

OpenStudy (anonymous):

Only negatives?

OpenStudy (ccswims):

so 10 in our case is our constant

OpenStudy (ccswims):

no

OpenStudy (ccswims):

just any number without a variable

OpenStudy (anonymous):

Oh, I remeber now. So it's Ax + By = C, x MUST be positive, so its 2x + 3y = -10, the 3y comes from the denominator, correct?

OpenStudy (ccswims):

2x + 3y = 10, its positive during the switch You get 3y because you multiplied -2/3 by 3, and you must do the same to the "y". Remember, the "y" was by itself and didn't have any number attached, but 3 times y makes it 3y

OpenStudy (anonymous):

So if it's -3/7 it would be, 3x +7y = 10?

OpenStudy (ccswims):

Yes, you understand it's because you multiplied 7 right?

OpenStudy (ccswims):

It's very simple when "y" is by itself, but most of the time "y" won't be by itself

OpenStudy (ccswims):

Try doing 5/8x + 2y = 10 by yourself

OpenStudy (ccswims):

I'm here if you need help

OpenStudy (anonymous):

So like, when Y not going to be by itself? Or would this be above 8th grade? (Algebra 1)

OpenStudy (ccswims):

I'm in 8th grade Algebra 1 and 9/10 times y is not by itself

OpenStudy (ccswims):

idk if your teacher is different but it's better if you understand also when y is not by itself

OpenStudy (ccswims):

you will be better prepared for the EOC (or whatever test you take in your state. I lie in Florida and we take the EOC)

OpenStudy (ccswims):

live in Florida**

OpenStudy (anonymous):

EOG, End of Grade So, 9/10y? or 9/10*y

OpenStudy (anonymous):

I'm going to try one, Slope: -6, (-2,4) would y= -6x + 4 be Correct?

OpenStudy (ccswims):

yes

OpenStudy (ccswims):

If you don't understand anything else search "Khan Academy" on google, they have GREAT videos that go in great detail in basically everything that makes up Algebra 1

OpenStudy (ccswims):

I myself must use the website sometimes

OpenStudy (ccswims):

I have to go now, but if I were you I would really consider staying up a little later or waking up earlier tomorrow and looking up "Khan Academy" to study for your test

OpenStudy (anonymous):

Thank you, :)